# CAST-Phys: Contactless Affective States Through Physiological signals Database

### v1

@@ Abstract
Index Terms:
Multi-modal database, Emotion recognition, Remote Photoplethysmography, Physiological signals
@@ I Introduction
Emotions are a vital part of human life, deeply influencing decision-making, behavior, and social interactions. They guide our actions and shape how we connect with others, playing a key role in every aspect of our interactions and decisions. Recently, Affective Computing [1], the field focused on developing systems to recognize, interpret, and simulate human emotions, has gained significant attention due to its broad range of applications in areas such as education [2, 3, 4] or healthcare [5, 6].
@@
Emotion states are commonly inferred in the literature using different modalities, such as facial expressions, speech, body gestures, and physiological signals. While facial expressions are increasingly popular due to their intuitive nature [7, 8, 9], physiological signals, also known as peripheral signals, offer distinct advantages. These include their growing availability, thanks to the rise of wearable devices, robustness to external visual noise (e.g., illumination), and high fidelity, as physiological reactions are involuntary and difficult to manipulate or conceal, making them more reliable indicators of the underlying emotional states. Several studies have established a significant link between emotional states and the Autonomic Nervous System (ANS) [10, 11]. The Sympathetic Nervous System (SNS) and the Parasympathetic Nervous System (PNS), both components of the ANS, play a crucial role in regulating various physiological signals. These signals include electrocardiogram (ECG), blood-volume pulse (BVP), electroencephalogram (EEG), electrodermal activity (EDA), breath rhythm (BR), among others.
@@
While individual modalities such as facial expressions, speech, or physiological signals have been used to infer emotional states, integrating models that combine features from multiple modalities has proven to be more effective. These multi-modal approaches enhance emotion recognition accuracy and have shown promising results in both research contexts and practical applications [12, 13, 14]. To develop these affective models, several emotion recognition databases for human-computer interaction (HCI) have been proposed, often focusing on facial expressions but less frequently combining them with physiological data. Early multimodal databases, such as MAHNOB-HCI [12], used implicit affective tagging to label content based on emotional responses. The DEAP database [15] examined emotional brain activity alongside peripheral signals, while the DECAF database [16] used magnetoencephalography (MEG) sensors to decode physiological responses to affective content. Recent databases like ASCERTAIN [17] and AMIGOS [18] include visual and physiological data (EEG, ECG, EDA), focusing primarily on peripheral signals. Despite these advancements, none are designed for remote data acquisition through facial video. The UBFC-Phys [19] and MMSE [20] datasets are the only ones aimed at remote physiological extraction. However, UBFC-Phys focuses only on stress, limiting its emotional scope, while the MMSE is not publicly available.
@@
  • We present the CAST-Phys database, a novel realistic multi-modal emotion recognition dataset designed to explore affective state recognition using both contact and non-contact physiological measurements from facial videos.
@@ II-A Multi-modal physiological emotion recognition databases
In recent decades, several emotion recognition datasets have been developed to explore emotion recognition across various modalities. However, only a few have been specifically designed for multi-modal emotion recognition, integrating modalities such as facial expressions, physiological signals, audio, and others. Among these, the MAHNOB-HCI [12] and DEAP [15] datasets, developed in 2011, were among the first multi-modal datasets. Both datasets recorded facial videos synchronized with physiological signals during emotion elicitation using videos, capturing both discrete and continuous emotional self-assessments. These datasets have become benchmark references for evaluating new approaches and remain widely used nowadays. Following their release, other similar multi-modal datasets emerged, including RECOLA [21], DECAF [16], and ASCERTAIN [17].
Subsequently, more sophisticated datasets were created. In 2016, the Multimodal Spontaneous Emotion Corpus (MMSE) was introduced, later extended to the BP4D+ database. This dataset includes data from 140 participants, featuring synchronized 3D and 2D facial videos, thermal imaging, and physiological data sequences, along with meta-data such as facial features and FACS codes. A subset of this dataset has been successfully used for video-based heart rate estimation [22, 23, 24], leveraging its photoplethysmography content. However, its potential for remote physiological emotion recognition remains largely unexplored, offering opportunities for future research.
Researchers such as Miranda et al. [18] and Song et al. [25] have also introduced the AMIGOS and MPED datasets, respectively, both of which improve emotion elicitation techniques to evoke more genuine emotional responses while recording physiological signals similar to those in earlier datasets. Finally, the most recent dataset in this field is the UBFC-Phys dataset [19], created in 2021 to study the impact of social stress on both contact and remote physiological responses. Although not specifically an emotion recognition dataset, it is the first dataset designed for remote physiological signal recovery aimed at extracting arousal responses.
@@ II-B Camera-based physiological signal measurement
Since Takano et al. [26] and Verkruysse et al. [27] demonstrated the feasibility of remote HR measurement from facial videos, researchers have developed various methods for physiological data recovery. Techniques include Blind Source Separation [28, 29], Normalized Least Mean Squares [30], self-adaptive matrix completion [22], and approaches based on the skin optical reflection model to mitigate motion artifacts [31, 32]. More recently, deep learning-based methods [33, 34, 35, 36, 37] have outperformed traditional techniques. Hybrid methods combine traditional approaches with CNNs for feature extraction [38, 39], while end-to-end models directly predict rPPG signals from facial videos [40, 41]. Transformer-based models [42, 43, 44, 45] capture long-range spatiotemporal features but face efficiency challenges, whereas lightweight frameworks [46, 47] focus on computational optimization. Unsupervised methods enhance generalization using meta-learning [48, 49], contrastive learning [50, 51], and techniques addressing domain gaps [52, 53, 54]. Additionally, data augmentation strategies tackle biases related to motion [55], skin tone [56], and heart rate distribution [57].
@@ II-C Camera-based physiological emotion recognition
Recent studies have explored the integration of physiological signals and facial expressions for remote emotion recognition. Benezeth et al. [58] were among the pioneers in demonstrating the feasibility of assessing emotional states using remote heart rate variability (rHRV). They employed frequency features from rHRV to differentiate between neutral and arousal states. Similarly, other works aimed to recognize emotions through the extraction of remote photoplethysmography (rPPG) signals. Sabour et al. [59] used remote pulse variability (rPV) to classify emotions into three discrete categories, while Yu et al. [41] applied remote heart rate (HR) analysis on the MAHNOB-HCI database, also extracting rHRV.
Until now, these studies have primarily focused on rHRV to estimate emotional states. Ouzar et al. [60] were the first to propose a bi-modal approach, combining remote HRV with facial features to improve emotion classification. More recently, Li and Peng [61] and Tao et al. [62] introduced similar bi-modal methods, combining facial expressions and rPPG signals using attention mechanisms for emotion recognition, applied to the MAHNOB-HCI and DEAP datasets, respectively. Although recent progress has been made, none of the proposed methods employs a multi-modal approach that incorporates more than two modalities, as we investigate in this work. Additionally, most existing methods evaluate their approaches on outdated multi-modal datasets, such as MAHNOB-HCI or DEAP, which were not designed for remote physiological signal sensing. These data sets overlook important factors, such as video compression [63, 64, 37] and the use of contact PPG signal acquisition, which are crucial for rPPG recovery. For these reasons, we introduce a novel database for remote physiological emotion recognition that will be presented in the next section.
@@ III-B Emotion Elicitation
In this work, the continuous affective representation [65] was used to characterize the emotional responses elicited by each stimulus. The affect circumplex model is a 2D emotion representation defined by two independent axes: valence and arousal. Valence reflects the level of pleasure or displeasure, while arousal indicates the degree of alertness or activation. To cover the entire valence-arousal domain, we divided the model into nine regions, as shown in Figure 1. Based on this segmentation, we have defined two video stimuli that target each of the resulting quadrants, with the objective of creating a dataset that uniformly covers the valence and arousal space, as detailed in the next subsection.
Figure 1: Affective circumplex model divided into 9 regions. The x-axis represents the valence, and the y-axis represents the arousal. Figure is taken from Paltoglou and Thelwall [66].
@@ III-C Stimuli database
Various methods have been used to elicit emotions, including images [67], audio [68], videos clips [12], and, more recently, virtual reality (VR) environments [69]. For the CAST-Phys dataset, we chose videos as they are multisensory, engaging both visual and auditory senses, without occluding facial regions for remote physiological extraction, unlike VR elicitation. The elicitation dataset consists of 18 videos selected to cover the nine regions of the valence-arousal space defined in Figure 1. The videos follow the criteria in [70]: short duration (30-60 seconds), self-explanatory visual scenes, and each designed to evoke a single emotion. Additionally, videos with significant spoken content are in Spanish, aligning with the cultural and linguistic context of the participants.
The stimuli dataset was compiled from various sources [12], [71], including publicly available online videos. To ensure the selection of the most representative video clips for each emotion, we used a majority-voting strategy with four external annotators. Each annotator independently rated the valence and arousal levels evoked by each video. After completing their annotations, videos that received large agreement from the annotators were included in the elicitation dataset. This rigorous process ensured the videos selected elicited consistent emotional responses.
@@ III-D Self-report
As stated in Subsection III-E, immediately after each trial, participants were instructed to use the Self-Assessment Manikin (SAM) test [72] to self-report the emotions they experienced while watching each video. The SAM test utilizes a set of five images, as shown in Figure 2, to represent valence and arousal independently. The images range from $-2$ to $+2$, allowing participants to indicate their perceived emotional valence and level of arousal.
@@ III-E Experimental protocol and acquisition setup
During the experiment, we first collected demographic information from participants and obtained their consent through the informed consent form, following the protocol approved by the Ethical Review Board (ERB). Participants were then introduced to the self-report system and the bio-sensors used in the experiment, followed by an initial test recording. Once data collection began, two test videos were shown, and after each, participants were asked to provide valence and arousal annotations. This stage also allowed participants to ask any questions about the rating system, or the concepts of valence and arousal. The annotations from these two videos were discarded and not used in the analysis. Following this, participants viewed the 18 selected elicitation videos and rated them after each one. A 30-second neutral interval was included between each emotional trial to prevent one emotion from influencing the next. Figure 3 summarizes the workflow of the experimental sessions.
@@
On the other hand, to allow participants to watch the videos and self-report their emotions, a graphic user interface (GUI) was developed using Psychopy v2023.1.0 [73]. The GUI ran on a separate computer from the data acquisition system, requiring synchronization between both computers to track when the participants were watching the videos. A LED light [^1] https://thepihut.com/products/tri-color-usb-controlled-tower-light-with-buzzer, positioned behind the participant, was used to mark the presentation of each stimulus. Controlled by the software running the GUI, the LED blinked a green light when a video started and ended, indicating the timing of the stimulus in the video recording. Additionally, artificial light sources were used to ensure uniform lighting conditions for all participants. Figure 5 shows an example of the experimental setup.
@@ IV-A1 Physiological signals
From the three collected physiological signals, various features are extracted using the NeuroKit2 toolbox [74]. Before feature extraction, all physiological signals are standardized by subtracting the mean and dividing by the standard deviation.
For the PPG signal, we derive HR and HRV, which are widely studied due to their strong correlation with emotional responses. From HR and HRV, we extract temporal, spectral, and non-linear features. Temporal features provide insight into beat-to-beat RR intervals, while spectral features, such as the ratio between low and high-frequency bands, offer information about the balance between the SNS and PNS components, also known as sympathovagal balance [75]. Non-linear features, such as the Poincaré plot, reveal correlations between successive RR intervals. In total, we extract 14 features from the HR signal. For the EDA signal, which reflects the SNS activity during high-intensity situations, we mainly considered its two main components: tonic Skin Conductance Level (SCL) and phasic Skin Conductance Response (SCR). SCL represents a smooth baseline level, whereas SCR captures rapid reactions to external stimuli. From the EDA signal, we extract six features. For the respiration signal, slow breathing is associated with relaxation, whereas irregular rhythms or rapid variations correspond to more aroused emotional states. We extract four respiration features. Across the three physiological modalities, we obtain a total of 24 features.

### v2

@@ Abstract
The CAST-Phys dataset is publicly available for research purposes upon request at IEEE DataPort platform.
@@ I Introduction
Emotions are a vital part of human life, deeply influencing decision-making, behavior, and social interactions. They guide our actions and shape how we connect with others, playing a key role in every aspect of our interactions and decisions. Recently, Affective Computing [picard2000affective], the field focused on developing systems to recognize, interpret, and simulate human emotions, has gained significant attention due to its broad range of applications in areas such as education [shen2009affective, mejbri2022trends, rodriguez2021affective] or healthcare [Liu, altameem2020facial].
@@
Emotion states are commonly inferred in the literature using different modalities, such as facial expressions, speech, body gestures, and physiological signals. While facial expressions are increasingly popular due to their intuitive nature [tian2001recognizing, bargal2016emotion, chen2016facial], physiological signals, also known as peripheral signals, offer distinct advantages. These include their growing availability, thanks to the rise of wearable devices, robustness to external visual noise (e.g., illumination), and high fidelity, as physiological reactions are involuntary and difficult to manipulate or conceal, making them more reliable indicators of the underlying emotional states. Several studies have established a significant link between emotional states and the Autonomic Nervous System (ANS) [kreibig2010autonomic, levenson2014autonomic]. The Sympathetic Nervous System (SNS) and the Parasympathetic Nervous System (PNS), both components of the ANS, play a crucial role in regulating various physiological signals. These signals include electrocardiogram (ECG), blood-volume pulse (BVP), electroencephalogram (EEG), electrodermal activity (EDA), breath rhythm (BR), among others.
@@
While individual modalities such as facial expressions, speech, or physiological signals have been used to infer emotional states, integrating models that combine features from multiple modalities has proven to be more effective. These multi-modal approaches enhance emotion recognition accuracy and have shown promising results in both research contexts and practical applications [soleymani2011multimodal, jung2019utilizing, comas2020end]. To develop these affective models, several emotion recognition databases for human-computer interaction (HCI) have been proposed, often focusing on facial expressions but less frequently combining them with physiological data. Early multimodal databases, such as MAHNOB-HCI [soleymani2011multimodal], used implicit affective tagging to label content based on emotional responses. The DEAP database [koelstra2011deap] examined emotional brain activity alongside peripheral signals, while the DECAF database [abadi2015decaf] used magnetoencephalography (MEG) sensors to decode physiological responses to affective content. Recent databases like ASCERTAIN [subramanian2016ascertain] and AMIGOS [miranda2018amigos] include visual and physiological data (EEG, ECG, EDA), focusing primarily on peripheral signals. Despite these advancements, none are designed for remote data acquisition through facial video. The UBFC-Phys [sabour2021ubfc] and MMSE [zhang2016multimodal] datasets are the only ones aimed at remote physiological extraction. However, UBFC-Phys focuses only on stress, limiting its emotional scope, while the MMSE is not publicly available.
@@
  • We present the CAST-Phys database, a novel realistic multi-modal emotion recognition dataset designed to explore affective state recognition using both contact and non-contact physiological measurements from facial videos. The dataset is freely available [^1] https://ieee-dataport.org/documents/cast-phys for research purposes upon request.
@@ II-A Multi-modal physiological emotion recognition databases
In recent decades, several emotion recognition datasets have been developed to explore emotion recognition across various modalities. However, only a few have been specifically designed for multi-modal emotion recognition, integrating modalities such as facial expressions, physiological signals, audio, and others. Among these, the MAHNOB-HCI [soleymani2011multimodal] and DEAP [koelstra2011deap] datasets, developed in 2011, were among the first multi-modal datasets. Both datasets recorded facial videos synchronized with physiological signals during emotion elicitation using videos, capturing both discrete and continuous emotional self-assessments. These datasets have become benchmark references for evaluating new approaches and remain widely used nowadays. Following their release, other similar multi-modal datasets emerged, including RECOLA [ringeval2013introducing], DECAF [abadi2015decaf], and ASCERTAIN [subramanian2016ascertain].
Subsequently, more sophisticated datasets were created. In 2016, the Multimodal Spontaneous Emotion Corpus (MMSE) was introduced, later extended to the BP4D+ database. This dataset includes data from 140 participants, featuring synchronized 3D and 2D facial videos, thermal imaging, and physiological data sequences, along with meta-data such as facial features and FACS codes. A subset of this dataset has been successfully used for video-based heart rate estimation [tulyakov2016self, liu2020multi, revanur2021first], leveraging its photoplethysmography content. However, its potential for remote physiological emotion recognition remains largely unexplored, offering opportunities for future research.
Researchers such as Miranda et al. [miranda2018amigos] and Song et al. [song2019mped] have also introduced the AMIGOS and MPED datasets, respectively, both of which improve emotion elicitation techniques to evoke more genuine emotional responses while recording physiological signals similar to those in earlier datasets. While all previous datasets were recorded under laboratory conditions, limiting their ecological validity, the DAPPER dataset [shui2021dataset] overcomes this limitation by collecting ambulatory physiological signals through wearable sensors and emotional self-reports from participants during their daily life activities. The dataset presented in this work, CAST-Phys, is intended to facilitate the study of remote emotion recognition through facial analysis, with the long-term goal of enabling affect estimation in real-world scenarios, similar to DAPPER, but without the need for wearable sensors. Finally, the most recent dataset in this field is the UBFC-Phys dataset [sabour2021ubfc], created in 2021 to study the impact of social stress on both contact and remote physiological responses. Although not specifically an emotion recognition dataset, it is the first dataset designed for remote physiological signal recovery aimed at extracting arousal responses.
@@ II-B Camera-based physiological signal measurement
Since Takano et al. [takano2007heart] and Verkruysse et al. [verkruysse2008remote] demonstrated the feasibility of remote HR measurement from facial videos, researchers have developed various methods for physiological data recovery. Techniques include Blind Source Separation [poh2010non, poh2010advancements], Normalized Least Mean Squares [li2014remote], self-adaptive matrix completion [tulyakov2016self], and approaches based on the skin optical reflection model to mitigate motion artifacts [de2013robust, wang2016algorithmic]. More recently, deep learning-based methods [vspetlik2018visual, yu2019remote, perepelkina2020hearttrack, lu2021dual, comas2024deep] have outperformed traditional techniques. Hybrid methods combine traditional approaches with CNNs for feature extraction [niu2019rhythmnet, song2021pulsegan], while end-to-end models directly predict rPPG signals from facial videos [chen2018deepphys, Yu2019RemotePS]. Transformer-based models [yu2023physformer++, gupta2023radiant, liu2024rppg, comas2025pulseformer] capture long-range spatiotemporal features but face efficiency challenges, whereas lightweight frameworks [liu2023efficientphys, comas2022efficient] focus on computational optimization. Unsupervised methods enhance generalization using meta-learning [lee2020meta, liu2021metaphys], contrastive learning [gideon2021way, sun2022contrast], and techniques addressing domain gaps [lu2023neuron, du2023dual, chari2024implicit]. Additionally, data augmentation strategies tackle biases related to motion [paruchuri2024motion, comas2025beatformer], skin tone [comas2024physflow], and heart rate distribution [hsieh2022augmentation].
@@ II-C Camera-based physiological emotion recognition
Recent studies have explored the integration of physiological signals and facial expressions for remote emotion recognition. Benezeth et al. [benezeth2018remote] were among the pioneers in demonstrating the feasibility of assessing emotional states using remote heart rate variability (rHRV). They employed frequency features from rHRV to differentiate between neutral and arousal states. Similarly, other works aimed to recognize emotions through the extraction of remote photoplethysmography (rPPG) signals. Sabour et al. [sabour2019emotional] used remote pulse variability (rPV) to classify emotions into three discrete categories, while Yu et al. [Yu2019RemotePS] applied remote heart rate (HR) analysis on the MAHNOB-HCI database, also extracting rHRV.
Until now, these studies have primarily focused on rHRV to estimate emotional states. Ouzar et al. [ouzar2022video] were the first to propose a bi-modal approach, combining remote HRV with facial features to improve emotion classification. More recently, Li and Peng [li2024end] and Tao et al. [tao2024facial] introduced similar bi-modal methods, combining facial expressions and rPPG signals using attention mechanisms for emotion recognition, applied to the MAHNOB-HCI and DEAP datasets, respectively. Although recent progress has been made, none of the proposed methods employs a multi-modal approach that incorporates more than two modalities, as we investigate in this work. Additionally, most existing methods evaluate their approaches on outdated multi-modal datasets, such as MAHNOB-HCI or DEAP, which were not designed for remote physiological signal sensing. These data sets overlook important factors, such as video compression [mcduff2017impact, nowara2021systematic, comas2024deep] and the use of contact PPG signal acquisition, which are crucial for rPPG recovery. For these reasons, we introduce a novel database for remote physiological emotion recognition that will be presented in the next section.
@@ III-B Emotion Elicitation
In this work, the continuous affective representation [russell1980circumplex] was used to characterize the emotional responses elicited by each stimulus. The affect circumplex model is a 2D emotion representation defined by two independent axes: valence and arousal. Valence reflects the level of pleasure or displeasure, while arousal indicates the degree of alertness or activation. To cover the entire valence-arousal domain, we divided the model into nine regions, as shown in Figure 1. Based on this segmentation, we have defined two video stimuli that target each of the resulting quadrants, with the objective of creating a dataset that uniformly covers the valence and arousal space, as detailed in the next subsection.
Figure 1: Affective circumplex model divided into 9 regions. The x-axis represents the valence, and the y-axis represents the arousal. Figure is taken from Paltoglou and Thelwall [paltoglou2012seeing].
@@ III-C Stimuli database
Various methods have been used to elicit emotions, including images [lang1997international], audio [nardelli2015recognizing], videos clips [soleymani2011multimodal], and, more recently, virtual reality (VR) environments [somarathna2022virtual]. For the CAST-Phys dataset, we chose videos as they are multisensory, engaging both visual and auditory senses, without occluding facial regions for remote physiological extraction, unlike VR elicitation. The elicitation dataset consists of 18 videos selected to cover the nine regions of the valence-arousal space defined in Figure 1. The videos follow the criteria in [nasoz2004emotion]: short duration (30-60 seconds), self-explanatory visual scenes, and each designed to evoke a single emotion. Additionally, videos with significant spoken content are in Spanish, aligning with the cultural and linguistic context of the participants.
The stimuli dataset was compiled from various sources [soleymani2011multimodal], [abadi2015decaf], [hadar2017implicit], including publicly available online videos. A total of 56 candidate videos were initially considered to select the final elicitation set. To ensure the selection of the most representative video clips for each emotion, we used a majority-voting strategy with four external annotators. Each annotator independently rated the valence and arousal levels evoked by each video. After completing their annotations, videos that received full agreement among the annotators were included in the elicitation dataset. The selection was conducted in five iterative rounds, where non-consensual videos were progressively replaced by new candidates until full agreement was achieved across all circumplex model regions. This rigorous process ensured the videos selected elicited consistent emotional responses. The final list of elicitation video clips used in the CAST-Phys dataset is presented in the supplementary material.
@@ III-D Self-report
As stated in Subsection III-E, immediately after each trial, participants were instructed to use the Self-Assessment Manikin (SAM) test [bradley1994measuring] to self-report the emotions they experienced while watching each video. The SAM test utilizes a set of five images, as shown in Figure 2, to represent valence and arousal independently. The images range from 2-2 to +2+2, allowing participants to indicate their perceived emotional valence and level of arousal.
@@ III-E Experimental protocol and acquisition setup
During the experiment, we first collected demographic information from participants and obtained their consent through the informed consent form, following the protocol approved by the Institutional Committee for Ethical Review of Projects (CIREP) at Pompeu Fabra University, application No. 306. Participants were informed that the experimental results could be published in academic journals or used for research purposes, asking them for permission to share anonymized raw data.
After providing consent, participants were then introduced to the self-report system and the bio-sensors used in the experiment, followed by an initial test recording. Once data collection began, two test videos were shown, and after each, participants were asked to provide valence and arousal annotations. This stage also allowed participants to ask any questions about the rating system, or the concepts of valence and arousal. The annotations from these two videos were discarded and not used in the analysis. Following this, participants viewed the 18 selected elicitation videos and rated them after each one. A 30-second neutral interval was included between each emotional trial to prevent one emotion from influencing the next. Figure 3 summarizes the workflow of the experimental sessions.
@@
The synchronization among the four modalities is essential for remote physiological signal sensing. To verify the alignment between the facial and physiological data, we computed the temporal offset, which showed an average difference of $8.19\pm 4.36$ milliseconds for the whole dataset, enabling precise temporal alignment for multimodal data acquisition.
On the other hand, to allow participants to watch the videos and self-report their emotions, a graphic user interface (GUI) was developed using Psychopy v2023.1.0 [peirce2019psychopy2]. The GUI ran on a separate computer from the data acquisition system, requiring synchronization between both computers to track when the participants were watching the videos. A LED light [^2] https://thepihut.com/products/tri-color-usb-controlled-tower-light-with-buzzer, positioned behind the participant, was used to mark the presentation of each stimulus. Controlled by the software running the GUI, the LED blinked a green light when a video started and ended, indicating the timing of the stimulus in the video recording.
To assess the accuracy of this synchronization, we extracted timestamps corresponding to the first and last LED flashes for every participant and trial. These timestamps were used to compute the effective stimulus duration and compared against the nominal duration of the selected elicitation videos. Across 1098 trial samples, the analysis revealed a Mean Absolute Error (MAE) of $17.10\pm 27.32$ ms. Given the monitor refresh rate and camera recording frequency of 60 Hz (frame duration $\approx$ 16.67 ms), this deviation confirms that synchronization precision is limited primarily by frame quantization rather than software drift. This demonstrates that the system achieves frame-level accuracy, enabling precise temporal alignment between the stimulus and the recorded videos.
Additionally, artificial light sources were used to ensure uniform lighting conditions for all participants. Figure 5 shows an example of the experimental setup.
@@ IV-A Data processing
The experimental validation was conducted using a total of 1098 trials, corresponding to 18 elicitation videos per participant across 61 participants. Owing to the accurate acquisition and the high quality of the recorded data, no trials were excluded, and all samples were included in the subsequent analyses.
@@ IV-A1 Physiological signals
From the three collected physiological signals, various features are extracted using the NeuroKit2 toolbox [Makowski2021neurokit]. Before feature extraction, all physiological signals are standardized by subtracting the mean and dividing by the standard deviation.
For the PPG signal, we derive HR and HRV, which are widely studied due to their strong correlation with emotional responses. From HR and HRV, we extract temporal, spectral, and non-linear features. Temporal features provide insight into beat-to-beat RR intervals, while spectral features, such as the ratio between low and high-frequency bands, offer information about the balance between the SNS and PNS components, also known as sympathovagal balance [Hori1992heart]. Non-linear features, such as the Poincaré plot, reveal correlations between successive RR intervals. In total, we extract 14 features from the HR signal. For the EDA signal, which reflects the SNS activity during high-intensity situations, we mainly considered its two main components: tonic Skin Conductance Level (SCL) and phasic Skin Conductance Response (SCR). SCL represents a smooth baseline level, whereas SCR captures rapid reactions to external stimuli. From the EDA signal, we extract six features. For the respiration signal, slow breathing is associated with relaxation, whereas irregular rhythms or rapid variations correspond to more aroused emotional states. We extract four respiration features. Across the three physiological modalities, we obtain a total of 24 features.